suppose that $f:[0,1]\rightarrow r$ is continous on [0,1] and differentiable on (0,1) with f(0)=0 anf f(1)=10.prove that there exist 100 distinct points x_k belonging to (0,1) such that summation from k=1 to 100 of 1/f'(x k)=1000


1
Expert's answer
2021-12-07T11:29:08-0500

equation of normal to f(x) at point (xk, yk):

"y-y_k=-(x-x_k)\/f'(x_k)"

so,

"\frac{1}{f(x_k)}=\frac{y-y_k}{x_k-x}"


then for each xk we can take values of x and y such that


"\displaystyle\sum_{k=1}^{100}\frac{1}{f(x_k)}=\displaystyle\sum_{k=1}^{100}\frac{y-y_k}{x_k-x}=1000"